An object that has mass of 28.0 kg is pushed along a horizontal surface with a force of 78.0N. if the force of friction is 65.0N, what is the magnitude of the acceleration of the object?


1
Expert's answer
2021-04-23T11:09:41-0400
"ma=F-F_f\\28a=78-65\\a=0.464\frac{m}{s^2}"